# Graph-Based Fault Diagnosis for Rotating Machinery: Adaptive Segmentation and Structural Feature Integration

### v1

@@ Abstract
This paper proposes a novel graph-based framework for robust and interpretable multiclass fault diagnosis in rotating machinery. The method integrates entropy-optimized signal segmentation, time–frequency feature extraction, and graph-theoretic modeling to transform vibration signals into structured representations suitable for classification. Graph metrics—such as average shortest path length, modularity, and spectral gap—are computed and combined with local features to capture both global and segment-level fault characteristics. Evaluated on two benchmark datasets, the CWRU bearing dataset (under 0–3 HP loads) and the SU gearbox and bearing datasets (under different speed–load configurations), the proposed method achieves high diagnostic accuracy. Classification scores reach up to 99.8% accuracy on Case Western Reserve University (CWRU) and 100% accuracy on the Southeast University datasets using a logistic regression classifier. Furthermore, the model exhibits strong noise resilience, maintaining over 95.4% accuracy at high noise levels (standard deviation = 0.5), and demonstrates excellent cross-domain transferability with up to 99.7% F1-score in load-transfer scenarios. Compared to traditional techniques, this approach requires no deep learning architecture, enabling lower complexity while ensuring interpretability. The results confirm the method’s scalability, reliability, and potential for real-time deployment in industrial diagnostics.
@@
Bearing fault diagnosis, entropy-based segmentation, feature noise robustness, graph-based learning, graph-theoretic features, multiclass classification, transfer learning, vibration signal analysis.
@@ 1 Introduction
Rotating machinery, such as motors, gearboxes, and bearings, is fundamental to industries including manufacturing, energy, transportation, and aerospace. The reliability of these systems depends on early fault detection in critical components like bearings and gearboxes, which are highly susceptible to failure due to mechanical stress (Lei et al., 2013). Undetected faults can cause costly downtime, maintenance expenses, or safety hazards, necessitating accurate and efficient fault diagnosis systems (Jardine et al., 2006).
Traditional fault diagnosis methods rely on statistical and spectral analysis of vibration signals, using fixed-window segmentation, handcrafted feature extraction (e.g., time- or frequency-domain metrics), and classifiers like support vector machines (Widodo and Yang, 2007). While effective in controlled settings, these approaches have limitations. Fixed segmentation struggles to adapt to dynamic signal variations across operating conditions, leading to suboptimal feature representations (Rauber et al., 2014). Handcrafted features often fail to capture complex or subtle fault patterns, especially in noisy environments (Lei et al., 2020). Additionally, these methods assume consistent training and testing conditions, which is rare in practice, resulting in poor generalization across varying loads, fault severities, or noise levels (Zhang et al., 2017).
Recent advances in deep learning have enhanced fault diagnosis capabilities. Convolutional neural networks (CNNs) (Wen et al., 2017), recurrent neural networks (RNNs) (Zhu et al., 2023), and attention-based models (Zhang et al., 2020a) have achieved strong results on benchmark datasets like the Case Western Reserve University (CWRU) bearing dataset (Center, 2014) and the Southeast University (SU) gearbox dataset(Wu, 2020). However, these methods require large labeled datasets, incur high computational costs, and lack interpretability due to their black-box nature (Zhang et al., 2020b). Their performance also degrades in cross-domain or cross-load scenarios, where training and deployment conditions differ (Li et al., 2022).
A critical gap in existing approaches is the underuse of structural information within vibration signals. These signals often exhibit temporal relationships reflecting physical dependencies, yet conventional methods process segments independently, overlooking inter-segment patterns (Yang et al., 2021). Graph-based representations offer a promising solution, enabling the modeling of structural relationships to enhance fault detection.
@@
To address these challenges, this paper proposes a novel graph-based framework for multiclass fault diagnosis in rotating machinery. Vibration signals are segmented adaptively using entropy-based optimization of window and step sizes, ensuring robust feature capture across diverse operating conditions. Each segment is represented by a comprehensive feature set, including time-domain statistics, frequency-domain metrics (e.g., spectral density, envelope analysis), and Teager-Kaiser energy (Welch, 1967). These features are clustered to form compact representations, from which k-nearest neighbor (kNN) graphs are constructed. Graph-theoretic metrics—such as average shortest path length, modularity, and spectral gap—encode fault-induced structural variations, capturing disruptions in signal regularity and community structure (Newman, 2018).
The framework integrates global graph metrics with local signal features, enabling a dual-level representation that distinguishes subtle fault types and enhances noise robustness. A multiclass logistic regression classifier, leveraging these high-quality features, achieves state-of-the-art performance with computational efficiency and interpretability, unlike complex deep learning models.
The objectives of this study are as follows:
@@
    To develop an adaptive segmentation technique based on entropy-driven optimization.
@@
    To construct graph representations from feature-rich segments of vibration signals.
@@
    To extract graph-theoretic metrics that effectively capture fault-related characteristics.
@@
    To evaluate the proposed framework on the CWRU and SU datasets across varying load conditions, fault types, noise levels, and cross-domain (transfer learning) scenarios.
Extensive experiments demonstrate that the proposed method achieves high accuracy, robust generalization, and resilience to noise, outperforming traditional and deep learning approaches. Its efficiency and interpretability make it well-suited for industrial applications. These results highlight the potential of graph-based learning to advance intelligent fault diagnosis, offering a scalable and practical solution for real-world challenges.
2 Literature Review
Traditional fault diagnosis methods rely on handcrafted features extracted from vibration signals, typically in the time or frequency domain, followed by classifiers like support vector machines (SVMs) (Widodo and Yang, 2007). Time-domain features (e.g., mean, RMS, skewness (Joanes and Gill, 1998)) and frequency-domain features (e.g., FFT (Welch, 1967), spectral envelope (Zheng et al., 2022)) capture fault signatures, while techniques like empirical mode decomposition (EMD) (Lei et al., 2013) and Teager-Kaiser energy (Kaiser, 1990) enhance signal decomposition. These methods achieve high accuracy in controlled settings, such as 95–98% on the CWRU dataset (Zhang et al., 2015), but struggle with dynamic operating conditions and noise (Rauber et al., 2014). Fixed segmentation and manual feature selection limit adaptability, requiring domain expertise (Lei et al., 2020).
Machine learning has improved fault diagnosis by automating feature selection and classification. Ensemble methods (Zhang et al., 2015) and transfer learning (Zhang et al., 2017) enhance robustness across operating conditions. For instance, Zhang et al. (Zhang et al., 2017) used neural networks for bearing fault diagnosis under varying loads, achieving 90–95% accuracy on CWRU cross-load scenarios. However, these methods still rely on handcrafted features, limiting their ability to capture complex patterns in noisy or high-dimensional data (Lei et al., 2020). Classifiers like k-nearest neighbors (kNN) (Cover and Hart, 1967) and SVMs (Widodo and Yang, 2007) are computationally efficient but lack the generalization of deep learning for large datasets.
Deep learning has revolutionized fault diagnosis by learning hierarchical features from raw vibration signals (Zhao et al., 2019). Convolutional neural networks (CNNs) (Wen et al., 2017; Niyongabo et al., 2022; Xu et al., 2025), long short-term memory (LSTM) networks (Zhang et al., 2020a), and attention-based models (Li et al., 2024) achieve high accuracies (98–99.9%) on benchmarks like CWRU and SU (Center, 2014; Wu, 2020). For example, Shao et al. (Shao et al., 2018) reported 99.5% accuracy using deep transfer learning on SU. Recent advancements include multi-source fusion (Makrouf et al., 2023), sparse CNNs (Xu et al., 2025), and transformer-based models (Zhou et al., 2024), improving efficiency and robustness. However, deep learning models require large labeled datasets, incur high computational costs (e.g., 536 s inference for DL-CNN (Shao et al., 2018)), and lack interpretability (Zhang et al., 2020b). Performance drops in cross-domain scenarios (e.g., 90–95% F1-score on cross-load CWRU (Li et al., 2022)) highlight generalization challenges (Zhu et al., 2023).
Graph-based methods offer a promising approach for modeling structural relationships in vibration signals, advancing fault diagnosis in rotating machinery (Yang et al., 2021; Singh, 2025). Yang et al. (Yang et al., 2021) employed spatial-temporal graphs to extract features, achieving 97% accuracy on the CWRU dataset, while Singh (Singh, 2025) integrated graph autoencoders with transformers for robust fault detection. These methods leverage graph-theoretic metrics, such as modularity (Newman, 2006), spectral gap (Cheeger, 1970; Chung, 1996), and small-world properties (Watts and Strogatz, 1998), to capture fault-induced disruptions in signal structure. While effective at modeling inter-segment dependencies, they are computationally intensive and limited in extracting frequency-domain features, as they prioritize structural relationships over spectral analysis (Yang et al., 2021; Lei et al., 2020). Few studies combine local signal characteristics with global graph metrics, a gap addressed by the proposed framework, which integrates both for enhanced accuracy and interpretability.
The CWRU bearing dataset (Center, 2014) and SU gearbox dataset (Wu, 2020) are widely used for benchmarking. CWRU includes vibration data under varying loads (0–3 HP) and fault types, while SU covers gearbox faults under diverse conditions (Shao et al., 2018). Performance metrics like accuracy, F1-score, and inference time are common, with statistical tests (e.g., ANOVA (Fisher, 1925), Tukey HSD (Tukey, 1949)) used to evaluate cross-load robustness (Zhang et al., 2017). Recent studies report accuracies of 98–99.9% on CWRU and SU, but cross-load performance remains a challenge (e.g., 90–95% F1-score for deep learning (Li et al., 2022)).
Despite significant advancements in vibration-based fault diagnosis of rotating machinery, several critical gaps and challenges persist across traditional, machine learning, deep learning, and graph-based approaches. Traditional and machine learning methods, reliant on manual feature engineering, lack adaptability to dynamic operating conditions and require extensive domain expertise to design effective features, limiting their performance in noisy or variable environments (Lei et al., 2020; Rauber et al., 2014). Deep learning models, while capable of automatically learning complex patterns, demand large labeled datasets and incur high computational costs, often rendering them impractical for real-time industrial applications (Zhu et al., 2023; Shao et al., 2018). Moreover, their black-box nature hampers interpretability, with post-hoc methods like Grad-CAM providing limited insight into fault mechanisms (Zhang et al., 2020b). Both traditional and deep learning approaches struggle with cross-domain generalization, as performance degrades significantly in cross-load or cross-domain scenarios due to domain shifts (Zhang et al., 2017; Li et al., 2022). Graph-based methods, although promising for modeling structural relationships in vibration signals, are underutilized and computationally intensive, often failing to integrate frequency-domain insights critical for fault detection (Yang et al., 2021; Singh, 2025). Additionally, the underuse of inter-segment relationships in vibration signals across all methods misses opportunities to capture fault-induced structural patterns, highlighting the need for approaches that combine adaptive feature extraction, structural modeling, and robust generalization (Lei et al., 2020; Yang et al., 2021).
@@ 3 Methodology
We propose a methodology that begins by extracting time- and frequency-domain features $\mathbf{F}_{i}$ from a multichannel vibrational signal $\mathbf{S}$. The signal is segmented, and clustered segments are used to construct a K-Nearest Neighbors (KNN) graph $\mathcal{G}$. From this graph, we compute key graph-theoretic features, including average shortest path length $L_{\text{avg}}$, modularity $Q_{\text{mod}}$, and spectral gap $\Delta_{\text{spec}}$. These features serve as input to a multiclass logistic regression model designed for fault detection. An overview of the proposed pipeline is illustrated in Figure 1, outlining the progression from raw signal input to fault classification via feature extraction, graph construction, graph-based feature computation, and model training. In the following subsection, we detail each component of the proposed approach.
Figure 1: Overview of the methodology for multiclass fault detection.
@@
Table 1: Summary of Key Symbols
@@
^$\hat{\ell}_{v}$ Predicted class label for node $v$
@@ 3.1 Optimizing Segmentation Process
$\beta=\lceil 2\cdot\text{IQR}(\mathbf{x}_{w})\cdot\omega^{-1/3}/(\max(\mathbf{x}_{w})-\min(\mathbf{x}_{w}))\rceil$
@@
    $p_{i}=S_{xx}(f_{i})/\left(\sum_{j=1}^{M}S_{xx}(f_{j})+\epsilon\right)$
@@
    $p_{i}=S_{\text{env}}(f_{i})/\left(\sum_{j=1}^{M}S_{\text{env}}(f_{j})+\epsilon\right)$

### v2

@@ Abstract
Rotating machinery, such as bearings and gearboxes, plays a critical role in industrial operations; however, faults in these components can lead to costly downtime, safety hazards, and inefficient maintenance. Traditional fault diagnosis methods often rely on manual expertise or complex deep learning models that require vast labeled data, high computational resources, and lack transparency, limiting their practicality in real-world settings. To address these challenges, this study introduces a novel graph-based framework for multiclass fault diagnosis that emphasizes interpretability, efficiency, and robustness. Vibration signals were adaptively segmented using entropy optimization to focus on informative fault-related bursts, and then transformed into time-frequency representations. Each segment was modeled as a graph, extracting structural features such as average shortest path length, modularity, and spectral gap, which were combined with statistical signal descriptors to train simple, interpretable classifiers: logistic regression, support vector machines, and random forests. Evaluated on the Case Western Reserve University (CWRU) bearing dataset across 0–3 HP loads and the Southeast University (SU) gearbox dataset under varied speed-load conditions, the framework outperformed traditional and deep learning approaches. The random forest classifier achieved up to 99.9% accuracy on the CWRU and 100% on the SU datasets. Even with added noise (standard deviation = 0.5), it maintains 89.3% and 98.1% accuracy on the CWRU and SU datasets, respectively. In cross-domain scenarios, it delivered F1-scores of 97.73% for cross-load and 97.99% for cross-fault transfers, showcasing strong generalizability without deep architectures. By combining graph-theoretic analysis with statistical signal processing, this method eliminates the need for resource-intensive models while enhancing adaptability to noise, varying loads, and diverse fault types. Its transparency allows engineers to trace anomalies to specific causes, making it ideal for real-time industrial fault diagnosis and predictive maintenance applications.
@@
Fault Diagnosis, Rotating Machinery, Graph-Based Framework, Vibration Signal Analysis, Random Forest Classifier
@@ 1 Introduction
Rotating machinery, including motors, gearboxes, and bearings, is integral to industries such as manufacturing, energy, transportation, and aerospace. The reliability of these systems hinges on early fault detection in critical components, which are prone to failure under mechanical stress (Lei et al., 2013). Undetected faults can lead to costly downtime, maintenance expenses, or safety hazards, necessitating accurate and efficient fault diagnosis systems (Jardine et al., 2006).Vibration signals, which capture fault-induced patterns, are widely used for diagnosis, but their analysis poses challenges due to dynamic operating conditions, noise, and complex fault patterns (Lei et al., 2020).
Traditional fault diagnosis methods for rotating machinery rely on handcrafted features extracted from vibration signals in the time or frequency domain, followed by classifiers like support vector machines (SVMs) (Widodo and Yang, 2007). Time-domain features (e.g., mean, RMS, skewness (Joanes and Gill, 1998)) and frequency-domain features (e.g., FFT (Welch, 1967), spectral envelope (Zheng et al., 2022)) capture fault signatures, while techniques like empirical mode decomposition (EMD) (Lei et al., 2013) and Teager-Kaiser energy (Kaiser, 1990) enhance signal decomposition. These methods achieve high accuracy (e.g., 95–98% on the CWRU dataset (Zhang et al., 2015)) in controlled settings but struggle with dynamic conditions and noise due to fixed segmentation and manual feature selection, requiring significant domain expertise (Lei et al., 2020; Rauber et al., 2014).
To address the limitations of manual feature engineering, machine learning methods automate feature selection and classification, marking a significant evolution in fault diagnosis (R and Mutra, 2025). Ensemble methods (Zhang et al., 2015) and transfer learning (Zhang et al., 2017) improve robustness across operating conditions. For example, Zhang et al. (Zhang et al., 2017) used neural networks for bearing fault diagnosis under varying loads, achieving 90-95% accuracy on CWRU cross-load scenarios. However, reliance on handcrafted features limits their ability to capture complex patterns in noisy or high-dimensional data. Classifiers like k-nearest neighbors (kNN) (Cover and Hart, 1967) and SVMs (Widodo and Yang, 2007; Yazdani-Asrami et al., 2023) are computationally efficient but lack the generalization of deep learning for large datasets (Lei et al., 2020).
Building on the automation introduced by machine learning, deep learning revolutionizes fault diagnosis by learning hierarchical features directly from raw vibration signals (Zhao et al., 2019). Convolutional neural networks (CNNs) (Wen et al., 2017; Niyongabo et al., 2022; Xu et al., 2025), long short-term memory (LSTM) networks (Zhang et al., 2020a), and attention-based models (Li et al., 2024) achieve high accuracies (98-99.9%) on benchmarks like CWRU and SU datasets (Center, 2014; Wu, 2020). For instance, Shao et al. (Shao et al., 2018) reported 99.5% accuracy using deep transfer learning on the SU dataset. Recent advancements include multi-source fusion (Makrouf et al., 2023), sparse CNNs (Xu et al., 2025), and transformer-based models (Zhou et al., 2024). However, deep learning models require large labeled datasets, incur high computational costs (e.g., 536 s inference for DL-CNN (Shao et al., 2018)), and lack interpretability, with performance dropping in cross-domain scenarios (e.g., 90-95% F1-score on cross-load CWRU (Li et al., 2022; Zhu et al., 2023)).
Several advanced methods have emerged to address the challenges of limited generalization and high computational costs in conventional deep learning approaches for fault diagnosis. For example, the MDIFN (Gao et al., 2024b) uses complex-valued CNNs and BiGRUs in an adversarial transfer-learning setup. This extracts domain-invariant spatiotemporal features from multi-source data, improving cross-domain performance and decreasing the need for extensive labeled datasets. However, its computational complexity and limited interpretability persist as drawbacks. Also, the Domain Feature Decoupling Network (DFDN) (Gao et al., 2024c) separates domain and fault features before adaptively merging them, increasing interpretability and data efficiency for cross-fault transfer learning.
While many traditional deep learning classifiers demand extensive labeled data, recent innovations mitigate this through lightweight designs and few-shot strategies. For example, Yan et al. (2024a) introduces LiConvFormer, a CNN/Transformer hybrid using separable multiscale convolution and broadcast self-attention for compact, robust fault diagnosis. Similarly, Yan et al. (2024b) proposes a few-shot class-incremental learning framework that enables wind turbine system-level fault diagnosis by incorporating new classes with minimal labeled examples, leveraging forward-backward compatible representations to avoid catastrophic forgetting. Additionally, Chen et al. (2024) develops a multi-modal self-supervised learning method that pre-trains on unlabeled vibration and acoustic signals for cross-domain one-shot bearing fault diagnosis, achieving high accuracy with just one target-domain label. These approaches collectively show that deep networks can deliver strong performance in data-scarce scenarios via efficient architectures and learning paradigms. Nevertheless, these techniques are compatible with the proposed framework, yielding improved accuracy and noise robustness, independent of deep learning, focusing on explainability and efficiency with limited data or domain changes.
Graph-based methods offer a novel approach by modeling structural relationships in vibration signals, addressing the adaptability issues of traditional and machine learning methods (Yang et al., 2021; Singh, 2025). Yang et al. (Yang et al., 2021) used spatial-temporal graphs to achieve 97% accuracy on CWRU, leveraging metrics like modularity (a measure of how well a graph divides into cohesive sub-groups (Newman, 2006)) and spectral gap (the difference between the largest eigenvalues of the graph’s adjacency matrix, indicating connectivity robustness (Cheeger, 1970; Chung, 1996)). Singh (Singh, 2025) integrated graph autoencoders with transformers for robust fault detection, using small-world properties (efficient connectivity with short average path lengths (Watts and Strogatz, 1998)) to capture fault-induced disruptions. These methods enhance adaptability to dynamic conditions by modeling inter-segment relationships, unlike the fixed segmentation of traditional approaches. However, they are computationally intensive and prioritize structural relationships over frequency-domain features (Yang et al., 2021; Lei et al., 2020). The proposed framework addresses this by integrating local signal characteristics with global graph metrics, improving both accuracy and interoperability.
The CWRU bearing dataset (Center, 2014) and SU gearbox dataset (Wu, 2020) are standard benchmarks. CWRU includes vibration data under varying loads (0–3 HP) and fault types, while SU covers gearbox faults under diverse conditions (Shao et al., 2018). Performance metrics like accuracy, F1-score, and inference time are evaluated, often using statistical tests (e.g., ANOVA (Fisher, 1925), Tukey HSD (Tukey, 1949)) for cross-load robustness (Zhang et al., 2017). Recent studies report 98-99.9% accuracy on these datasets, but cross-load performance remains lower (90-95% F1-score (Li et al., 2022)).
Despite advancements, challenges persist across fault diagnosis methods. Traditional and machine learning approaches lack adaptability to dynamic conditions due to manual feature engineering (Lei et al., 2020; Rauber et al., 2014). Deep learning models, while powerful, require large datasets, incur high computational costs, and lack interpretability, with post-hoc methods like Grad-CAM offering limited insight (Zhang et al., 2020b). Cross-domain generalization remains a hurdle, with performance drops due to domain shifts (Zhang et al., 2017; Li et al., 2022). Graph-based methods, though effective for structural modeling, are computationally intensive and underutilize frequency-domain features (Yang et al., 2021; Singh, 2025). The underuse of inter-segment relationships in vibration signals across all methods misses opportunities to capture fault-induced patterns, underscoring the need for adaptive, structurally informed, and generalizable approaches (Lei et al., 2020; Yang et al., 2021).
To address these challenges, this paper proposes a novel graph-based framework that emphasizes interpretability and computational efficiency through feature engineering for multiclass fault diagnosis in rotating machinery. Unlike end-to-end deep learning, our approach integrates entropy-driven segmentation to optimize window and step sizes, ensuring adaptive and robust feature extraction under varying conditions.
@@
Each segment is characterized by a comprehensive feature set, including time-domain statistics, frequency-domain descriptors (e.g., spectral density, envelope analysis), and Teager–Kaiser energy (Welch, 1967). These features are clustered to form representative patterns, from which $k$-nearest neighbor (kNN) graphs are constructed. Graph-theoretic metrics, such as average shortest path length, modularity, and spectral gap, capture fault-induced structural changes (Newman, 2018), providing a lightweight yet powerful basis for fault classification. By combining global graph metrics with local signal features, the framework achieves a dual-level representation that enhances the detection of subtle faults and improves noise resilience.
To leverage this enriched representation for classification, the proposed framework integrates a multiclass strategy using logistic regression (LR), random forests (RF), and support vector machines (SVM). These classifiers are chosen to balance interpretability, generalization, and robustness: LR provides a transparent baseline; SVM captures complex decision boundaries in high-dimensional space; and RF mitigates feature redundancy through ensemble learning. Integrating Shapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017) further improves interpretability by quantifying feature contributions, reducing model-specific bias, and promoting trustworthy fault explanations. The complete pipeline is validated on the benchmark CWRU and SU datasets across diverse load and fault conditions, demonstrating superior diagnostic accuracy, robustness to noise, and cross-domain adaptability compared to conventional, machine learning, and deep learning approaches.
This study aims to:
@@
    Enhance vibration signal interpretation through entropy-guided segmentation.
@@
    Leverage graph learning to capture structural patterns in machinery health data.
@@
    Combine graph-based and statistical features for robust fault characterization.
@@
    Evaluate scalable, interpretable classifiers on standard benchmark datasets.
  • 5.
    Demonstrate robustness to noise and transferability across domains.
This approach bridges the gap between structural modeling and feature engineering, offering a transparent, computationally efficient, and generalizable solution for real-time industrial monitoring and predictive maintenance.
@@ 2 Methodology
We propose a methodology that begins by extracting time- and frequency-domain features $\mathbf{F}_{i}$ from a multichannel vibrational signal $\mathbf{S}$. The signal is segmented, and clustered segments are used to construct a K-Nearest Neighbors (KNN) graph $\mathcal{G}$. From this graph, we compute key graph-theoretic features, including average shortest path length $L_{\text{avg}}$, modularity $Q_{\text{mod}}$, and spectral gap $\Delta_{\text{spec}}$. These features serve as input to a multiclass classifier model designed for fault detection. An overview of the proposed pipeline is illustrated in Figure 1, outlining the progression from raw signal input to fault classification via feature extraction, graph construction, graph-based feature computation, and model training. In the following subsection, we detail each component of the proposed approach.
Figure 1: Overview of the methodology for multiclass fault detection. Source: Authors own work
@@
Table 1: Summary of Key Symbols. Source: Authors own work.
@@
$\hat{\ell}_{v}$ Predicted class label for node $v$
@@ 2.1 Optimizing Segmentation Process
$\beta=\lceil 2\cdot\text{IQR}(\mathbf{x}_{w})\cdot\omega^{-1/3}/(\max(\mathbf{x}_{w})-\min(\mathbf{x}_{w}))\rceil$ (2)
@@
    $p_{i}=S_{xx}(f_{i})/\left(\sum_{j=1}^{M}S_{xx}(f_{j})+\epsilon\right)$ (8)
@@
    $p_{i}=S_{\text{env}}(f_{i})/\left(\sum_{j=1}^{M}S_{\text{env}}(f_{j})+\epsilon\right)$ (11)